OpenStudy (anonymous):

A car with a mass of 1.1 × 103 kilograms hits a stationary truck with a mass of 2.3 × 103 kilograms from the rear end. The initial velocity of the car is +22.0 meters/second. After the collision the velocity of the car is -11.0 meters/second. What is the velocity of the truck after this elastic collision?

OpenStudy (anonymous):

it is going to be conservation of momentum, and use the formula KE=mv^2 but I cant get the right answer! HELP!!!

OpenStudy (anonymous):

you need to use only conservation of momentum and you will get the answer make sure you use correct signs

OpenStudy (anonymous):

So is that formula right? I keep getting answers that aren't options

OpenStudy (anonymous):

+13 meters/second +20 meters/second +15.7 meters/second +7.5 meters/second +9.8 meters/second THOSE are my options :/

OpenStudy (anonymous):

solution: http://openstudy.com/study#/updates/535039a8e4b048ad9ac49fc9

OpenStudy (anonymous):

Hmm, momentum is (mass x velocity) initial momentum of car : = (1.1 x 10^3 kg) x (22.0 m/sec) = 2.42 x 10^4 kg•m / sec final momentum of car : = (1.1 x 10^3 kg) x (-11.0 m/sec) = -1.21 x 10^4 kg•m / sec so the car 'lost' the difference in energy or (2.42 x 10^4 - 1.21 x 10^4) = 1.21 x 10^4 kg•m / sec This energy was transferred to the stationary truck, which means that the truck's momentum becomes p = mv or 1.21 x 10^4 kg•m / sec = (2.3 x 10^3 kg) * v and rearranging that to solve for v gives v = (1.21 x 10^4 kg•m / sec) / (2.3 x 10^3 kg) ≃ + 5.3 m/s